OpenStudy (anonymous):

what does it mean cos7(root over 5)x i mean how it should be written

OpenStudy (anonymous):

you mean \[\cos(\sqrt[7]{5}x)\]? then\[=\cos\left(5^{1\over7}x\right)\]

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

were you asking how to write that in TeX?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\sqrt[n]{x} produces \[\sqrt[n]{x}\]